OpenStudy (anonymous):

how do you find the area of a triangle with B=4.2in. and height H=6.3??

OpenStudy (mathmale):

The formula for the area of a triangle of base b and height h is (1/2)bh. Try that.

OpenStudy (anonymous):

to find the area of a triangle u use the formula 1/2bxh=a so 4.2x6.2=24.06/0.5=52.08